OpenStudy (anonymous):

Which of these values of x satisfies the inequality −2x + 4 ≤ −6 A. 5 B. 0 C. −2 D. −6

OpenStudy (lynfran):

or we can solve for x we get \[-2x+4\le -6\]\[-2x \le -6-4\]\[-2x \le -10\]\[x \ge 5\]

OpenStudy (anonymous):

Ok thanks

OpenStudy (lynfran):

welcome